If 100 ml of water is added to 700 ml of a 1 in 600 solution of cetrimide, what is the strength of the resulting cetrimide solution expressed as a percentage?
Answer #1

Quantity 1= 700 ml of cetrimide

Concentration 1 = 1 in a 600 solution = 0.000016 %

Quantity 2= 100 ml of water

Concentration 2= X

Q1 * C1= Q2* C2

700* 0.000016= 100* X

X(C2)= 0.000112 = 0.0112 %